OpenStudy (anonymous):

how do I extract the square roots in 2(x-5)^2 = 17

OpenStudy (anonymous):

2(x^2+25-10x)=17 do it urself after it

OpenStudy (anonymous):

isn't that just factoring?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but that's not extracting the square :(

OpenStudy (anonymous):

and it doesn't factor nicely.

OpenStudy (anonymous):

2x^2-20x+8

OpenStudy (anonymous):

+8? 25(2) = 50 -17 = 33

OpenStudy (anonymous):

x^2-10x+4 you must rooting

OpenStudy (anonymous):

oh no you do this in two steps only

OpenStudy (anonymous):

\[2(x-5)^2=17\] \[(x-5)^2=\frac{17}{2}\] \[x-5=\pm\sqrt{\frac{17}{2}}\] \[x=5\pm\sqrt{\frac{17}{2}}\]